# Axial Neural Networks for Dimension-Free Foundation Models

### v1

@@ 3.3 Example: Dimension-Agnostic PDE Solver
We provide an example of a dimension-agnostic PDE solver by merging GXNN and MPP. MPP consists of patch embedding, multiple attention layers, and patch de-embedding. The patch embedding and patch de-embedding are CNNs. Thus, in the axial implementation, we use the patch embedding as the lifting layer and the rest as the subsequent layers. The attention layers are the same as the $\operatorname{SelfAttn}$ example described in Eq.˜19, and the patch de-embedding is a convolution variant of it. The details and illustrations of the example can be referred to in Appendix˜A.
@@ 5.1 Toy Dataset: Gaussian Process Kernel Prediction
Comparison Between SXNN and GXNN. SXNN exhibits a trade-off between computational efficiency and expressivity inefficiency as described in Eq.˜12. To examine this trade-off and verify the necessity of GXNN, we build a set-based axial CNN (SXCNN) composed of Conv1Ds and a graph-based axial CNN (GXCNN) from the previous experiment. We then compare their number of parameters (#Params), wall clock time (W. Clock) of forward computation, and inference performance (Acc.) using the GP toy dataset. Table˜1 shows the comparisons, where SXCNN-L is an enlarged version of SXCNN for fair comparison with GXCNN. Note that SXCNN with the same depth and width as GXCNN has $6\times$ fewer parameters and 80% reduced wall clock time, but its performance degrades in high dimensions. After increasing the depth and width, it performs fairly well in high dimensions, but GXCNN still exhibits superior performance, indicating the necessity of GXNN structures.
@@ 5.2 PDE Solver Foundation Models
We implemented our architecture in two state-of-the-art PDE foundation model baselines: CViT [35] and MPP [28]. Both are based on ViT [12], incorporating patch embedding and multiple attention layers with some revisions optimized for PDE learning. Note that these baselines are designed only for 2D data, as the 2D convolutional layer for patch embedding is dimension-dependent, though the attention layers are not. We build their XNN variants, termed X-CViT and X-MPP, which include axial linears, axial convolutions, and axial attentions. The architectural specifications are detailed in Appendix˜C.
@@
5.2.1 Expressivity of XNN
We evaluated the architectural expressivity of XNN by measuring how well it performs compared to dimension-specific models in three different settings: single PDE training, multiple PDE training, and single PDE finetuning. Since the baselines are designed for 2D PDEs, we evaluate performance only on 2D PDEs: Diffusion-Reaction (DR), Incompressible Navier-Stokes (NS), Shallow Water Equation (SWE), and Compressible Fluid Dynamics (CFD) with Mach numbers $M=0.1$ and $1.0$. We measure NRMSE for each PDE. Notably, for 1D-2D joint training in MPP, we convert 1D PDE solutions to 2D by padding with zero values.
@@ 5.2.1 Expressivity of XNN
5.2.2 Unseen Dimension Generalization
In this experiment, we demonstrate the dimension-generalization (few-shot learning) ability of XNN on data from unseen dimensions. To do so, we pretrain both MPP and X-MPP on 2D PDE data and compare their finetuning performance on unseen 1D and 3D PDEs, demonstrating the benefit of multi-dimensional pretraining in XNN-based foundation models. We use three 1D PDEs and one 3D PDE: Diffusion Sorption (DS), Burgers’ equation (BG), 1D Compressible Fluid Dynamics (CFD 1D), and 3D Compressible Fluid Dynamics (CFD 3D). We evaluate using the same metric, NRMSE, and compare how finetuning performance improves as the size of the given dataset increases.
@@ 6 Conclusion
7 Limitations and Future Work
An axial version of the cross attention between two tensors of different dimensions is not yet fully explored. Addressing this issue as a natural extension of this work is a valuable future direction that would further widen the usability of XNNs. Additionally, improving the computational efficiency of GXNN, especially for high-dimensional data, is also a promising direction towards the practical deployment of XNNs in large-scale scientific simulations.
@@ Acknowledgement
We thank Heejun Lee for his thoughtful technical support, Jules Berman for his insightful concerns, and Kyunghyun Cho for facilitating our wonderful collaborative research. This work was partly supported by Institute of Information & communications Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT) (No.RS-2019-II190075, Artificial Intelligence Graduate School Program(KAIST); No.RS-2024-00509279, Global AI Frontier Lab; No.RS-2022-II220713, Meta-learning Applicable to Real-world Problems). H. Yang was supported by the National Research Foundation of Korea(NRF) grant funded by the Korean Government(MSIT) (No. RS-2023-00279680).
@@ Appendix A Dimension-Agnostic PDE Solver
where $\operatorname{Norm}$ indicates InstanceNorm [34]. On the other hand, in the axial version, we use $\operatorname{Embedder}$ for $\psi$ in Eq.˜17 with $\max$ aggregation as follows:
$\begin{aligned} &\tilde{\boldsymbol{h}}_{d_{i}}=\max_{j\neq i,\ j\in[1,K]}\text{T}_{(K-1)(j)}\left(\operatorname{AvgPool}_{1,\ldots,K-2}\left(\operatorname{Embedder}_{K-1,K}\left(\text{T}_{(j)(K-1)}\left(\text{T}_{(i)(K)}(\boldsymbol{x})\right)\right)\right)\right),\\ &\boldsymbol{h}_{d_{i}}=\operatorname{Norm}(\tilde{\boldsymbol{h}}_{d_{i}}),\end{aligned}$ (24)
@@
In contrast, in the graph-based axial case, we obtain $K$ features due to the lifting layer in the patch embedding. Therefore, we also use a set-based structure but apply the same attention to each feature independently, which differs from the graph-based attention example in Eq.˜20. The self-attention used is:
@@
In the axial version, we define two $\operatorname{Deembedder}$ blocks: one for updating the node feature and the other for updating the neighborhood feature (nhbr). We aggregate them using a $\max$ operation. Then, Eq.˜21 becomes:
@@ Appendix B Proofs of Theorems
Let $\phi$ be an axis-permutation equivariant function (e.g., another SXNN or pointwise operation). The SXNN in Eq.˜8 is axis-permutation equivariant for any rank-$K$ tensor $\boldsymbol{x}\in\mathbb{R}^{d_{1}\!\times\!d_{2}\!\times\!\ldots\!\times\!d_{K}}$.
@@
We prove the axis-permutation equivariance property defined in Eq.˜6. For any axis permutation $\Pi^{\prime}\in\boldsymbol{\Pi}$, the permutation of the input in Eq.˜8 before applying $\phi$ becomes
@@
which proves that Eq.˜8 is axis-permutation equivariant for all ${\Pi^{\prime}}\in\boldsymbol{\Pi}$. ∎
@@
The lifting layer of GXNN, Eq.˜17, is axis-permutation equivariant for any rank-$K$ tensor $\boldsymbol{x}\in\mathbb{R}^{d_{1}\!\times\!d_{2}\!\times\!\ldots\!\times\!d_{K}}$, except when $K=1$.
Proof.
The lifting operation converts an axis permutation of the input into an element-wise permutation of the output features. Therefore, we prove permutation equivariance instead of axis-permutation equivariance. For the lifting operation $f$,
$f(d_{i})=\phi\bigg(d_{i},\bigoplus_{j\neq i}^{K}\psi(d_{i},d_{j})\bigg),$ (32)
which is applied to each axis in the same way. Under an axis permutation $\pi$, this becomes:
$\begin{bmatrix}f(\pi(d_{1}))\\ f(\pi(d_{2}))\\ \vdots\\ f(\pi(d_{K}))\end{bmatrix}=\begin{bmatrix}f(d_{\pi(1)})\\ f(d_{\pi(2)})\\ \vdots\\ f(d_{\pi(K)})\end{bmatrix}=\pi\begin{bmatrix}f(d_{1})\\ f(d_{2})\\ \vdots\\ f(d_{K})\end{bmatrix}$, (33)
which shows that a permutation of the input axes results in the same permutation of the output features. ∎
B.3 ˜3.3
The subsequent layers of GXNN, Eq.˜21, are axis-permutation equivariant for any rank-$K$ tensor $\boldsymbol{h}_{d_{i}}\in\mathbb{R}^{d_{1}\!\times\!d_{2}\!\times\!\ldots\!\times\!d_{K}}$.
Proof.
From Eq.˜33, we observe that
$\boldsymbol{h}_{d_{\pi(i)}}=f(d_{\pi(i)})=\pi(f(d_{i}))=\pi(\boldsymbol{h}_{d_{i}}).$ (34)
Using this, for the subsequent layer operation $g$ described in Eq.˜21, the permutation of the input becomes
$\begin{bmatrix}g(\pi(\boldsymbol{h}_{d_{1}}))\\ g(\pi(\boldsymbol{h}_{d_{2}}))\\ \vdots\\ g(\pi(\boldsymbol{h}_{d_{K}}))\end{bmatrix}=\begin{bmatrix}g(\boldsymbol{h}_{\pi(d_{1})})\\ g(\boldsymbol{h}_{\pi(d_{2})})\\ \vdots\\ g(\boldsymbol{h}_{\pi(d_{K})})\end{bmatrix}=\pi\begin{bmatrix}g(\boldsymbol{h}_{d_{1}})\\ g(\boldsymbol{h}_{d_{2}})\\ \vdots\\ g(\boldsymbol{h}_{d_{K}})\end{bmatrix}$. (35)
Note that the element-wise permutation equivariance is translated back to axis permutation equivariance through the pooling layer at the end. ∎
B.4 Examples in Eq.˜15, Eq.˜19, and Eq.˜20 in GXNN
The examples extend the constructions of ˜3.2 and ˜3.3, preserving the axis order information during all aggregations. However, when using axis alignment through $\Pi$ or T, as in Eq.˜15, Eq.˜19, and Eq.˜20, additional assumptions are required to ensure axis-permutation equivariance.
@@
With these assumptions, we prove that GXNN, consisting of the lifting, subsequent layers, and pooling, is axis-permutation equivariant as follows:
$\operatorname{Pooling}(\operatorname{Subsequent}(\operatorname{Lifting}(\Pi(\boldsymbol{x}))))=\Pi(\operatorname{Pooling}(\operatorname{Subsequent}(\operatorname{Lifting}(\boldsymbol{x})))),\quad\forall\Pi\in\boldsymbol{\Pi}.$ (36)
@@
The lifting layer $\operatorname{Lifting}$ described in Eq.˜15 can be equivalently written as
$\boldsymbol{h}_{d_{i}}=\phi\left(\text{T}_{(i)(K)}(\boldsymbol{x}),\sum_{j\neq i}^{K}\text{T}_{(K-1)(j)}\psi\left(\text{T}_{(j)(K-1)}\text{T}_{(i)(K)}(\boldsymbol{x})\right)\right),$ (37)
where $\text{T}_{(m)(n)}(\boldsymbol{x})$ denotes the transpose between the $m$-th and $n$-th axes. Applying an axis permutation $\Pi(\boldsymbol{x})$ to the input gives
$\phi\left(\text{T}_{(i)(K)}(\Pi(\boldsymbol{x})),\sum_{j\neq i}^{K}\text{T}_{(K-1)(j)}\psi\left(\text{T}_{(j)(K-1)}\text{T}_{(i)(K)}(\Pi(\boldsymbol{x}))\right)\right).$ (38)
Substituting $\Pi^{\prime}\text{T}_{(k)(K)}=\text{T}_{(i)(K)}\Pi$ and $i=\pi(k)$, where $\Pi^{\prime}$ permutes only axes $1$ through K1K-1, we get
$=\phi\left(\Pi^{\prime}\text{T}_{(k)(K)}(\boldsymbol{x}),\sum_{j\neq\pi(k)}^{K}\text{T}_{(K-1)(j)}\psi\left(\text{T}_{(j)(K-1)}\Pi^{\prime}\text{T}_{(k)(K)}(\boldsymbol{x})\right)\right).$ (39)
Let $\pi$ be the element-wise permutation across axes induced by $\Pi$. Next, substituting $\Pi^{\prime\prime}\text{T}_{(l)(K-1)}=\text{T}_{(j)(K-1)}\Pi^{\prime}$, where $\Pi^{\prime\prime}$ permutes only axes $1$ through K2K-2, we obtain
$=\phi\left(\Pi^{\prime}\text{T}_{(k)(K)}(\boldsymbol{x}),\sum_{\pi^{\prime}(l)\neq\pi(k)}^{K}\Pi^{\prime}\text{T}_{(l)(K-1)}{\Pi^{\prime\prime}}^{-1}\psi\left(\Pi^{\prime\prime}\text{T}_{(l)(K-1)}\text{T}_{(k)(K)}(\boldsymbol{x})\right)\right).$ (40)
Here, $\pi^{\prime}$ is the element-wise permutation associated with $\Pi^{\prime}$. Since the summation is permutation-invariant and $\pi^{\prime}(l)\neq\pi(k)$ is equivalent to $l\neq k$ (as $\pi^{\prime}$ does not act on the $K$-th axis and $k$-th axis is transposed to $K$ by $\text{T}_{(k)(K)}$), and using ˜2 (permutation equivariance of $\psi$ under $\Pi^{\prime\prime}$), we have
$=\phi\left(\Pi^{\prime}\text{T}_{(k)(K)}(\boldsymbol{x}),\sum_{l\neq k}^{K}\Pi^{\prime}\text{T}_{(l)(K-1)}{\Pi^{\prime\prime}}^{-1}\Pi^{\prime\prime}\psi\left(\text{T}_{(l)(K-1)}\text{T}_{(k)(K)}(\boldsymbol{x})\right)\right).$ (41)
Applying ˜2 again for $\phi$ under $\Pi^{\prime}$ yields
$=\Pi^{\prime}\phi\left(\text{T}_{(k)(K)}(\boldsymbol{x}),\sum_{l\neq k}^{K}\text{T}_{(l)(K-1)}\psi\left(\text{T}_{(l)(K-1)}\text{T}_{(k)(K)}(\boldsymbol{x})\right)\right)=\Pi^{\prime}\boldsymbol{h}_{d_{k}}=\Pi^{\prime}\boldsymbol{h}_{d_{\pi^{-1}(i)}}.$ (42)
Next, the subsequent layers $\operatorname{Subsequent}$ described in Eq.˜19 can be equivalently written as
$\boldsymbol{h}^{\prime}_{d_{i}}=\phi\left(\boldsymbol{h}_{d_{i}},\text{T}_{(i)(K)}\sum_{j=1}^{K}\text{T}_{(K)(j)}\psi(\boldsymbol{h}_{d_{j}})\right).$ (43)
Using Eq.˜42, permutation of input leads to
$\phi\left(\Pi^{\prime}\boldsymbol{h}_{d_{k}},\text{T}_{(\pi^{\prime}(k))(K)}\sum_{\pi^{\prime}(l)=1}^{K}\text{T}_{(K)(\pi^{\prime}(l))}\psi(\Pi^{\prime}\boldsymbol{h}_{d_{l}})\right).$ (44)
@@
$=\Pi^{\prime}\phi\left(\boldsymbol{h}_{d_{k}},\text{T}_{(k)(K)}\sum_{l=1}^{K}\text{T}_{(K)(l)}\psi(\boldsymbol{h}_{d_{l}})\right)=\Pi^{\prime}\boldsymbol{h}^{\prime}_{d_{k}}.$ (45)
Finally, the pooling layer aggregates the axis-wise features into a single feature, which remains equivariant to axis permutations $\Pi$. As described in the example in Eq.˜22, it is equivalently written as
$\boldsymbol{h}^{\prime\prime}=\sum_{j=1}^{K}\text{T}_{(K)(j)}\boldsymbol{h}^{\prime}_{d_{j}}.$ (46)

### v2

@@ 3.3 Example: Dimension-Agnostic PDE Solver
We provide an example of a dimension-agnostic PDE solver by merging GXNN and MPP. MPP consists of patch embedding, multiple attention layers, and patch de-embedding. The patch embedding and patch de-embedding are CNNs. Thus, in the axial implementation, we use the patch embedding as the lifting layer and the rest as the subsequent layers. The attention layers are the same as the $\operatorname{SelfAttn}$ example described in §˜3.2, and the patch de-embedding is a convolution variant of it. The details and illustrations of the example can be referred to in Appendix˜A.
@@ 5.1 Toy Dataset: Gaussian Process Kernel Prediction
Comparison Between SXNN and GXNN. SXNN exhibits a trade-off between computational efficiency and expressivity inefficiency as described in Eq.˜13. To examine this and verify the necessity of GXNN, we build a set-based axial CNN (SXCNN) composed of Conv1Ds and a graph-based axial CNN (GXCNN) from the previous experiment. We then compare their number of parameters (#Params), wall clock time (W. Clock) of forward computation, and inference performance (Acc.) using the GP toy dataset. Table˜1 shows the comparisons, where SXCNN-L is an enlarged version of SXCNN for fair comparison with GXCNN. Note that SXCNN with the same depth and width as GXCNN has $6\times$ fewer parameters and 80% reduced wall clock time, but its performance degrades in high dimensions. After increasing the depth and width, it performs fairly well in high dimensions, but GXCNN still exhibits superior performance, indicating the necessity of GXNN.
@@ 5.2 PDE Solver Foundation Models
We implemented our architecture in two state-of-the-art PDE foundation model baselines: CViT [35] and MPP [28]. Both are based on ViT [12], incorporating patch embedding and multiple attention layers with some revisions optimized for PDE learning. Note that these baselines are designed only for 2D data, as the 2D convolutional layer for patch embedding is dimension-dependent, though the attention layers are not. We build their XNN variants, termed X-CViT and X-MPP, which include axial linears, axial convolutions, and axial attentions. The specifications are detailed in Appendix˜C.
@@
Expressivity of XNN. We evaluated the architectural expressivity of XNN by measuring how well it performs compared to dimension-specific models in three different settings: single PDE training, multiple PDE training, and single PDE finetuning. Since the baselines are designed for 2D PDEs, we evaluate performance only on 2D PDEs: Diffusion-Reaction (DR), Incompressible Navier-Stokes (NS), Shallow Water Equation (SWE), and Compressible Fluid Dynamics (CFD) with Mach numbers $M=0.1$ and $1.0$. We measure NRMSE for each PDE. Notably, for 1D-2D joint training in MPP, we convert 1D PDE solutions to 2D by padding with zero values.
@@
Unseen Dimension Generalization. In this experiment, we demonstrate the dimension-generalization (few-shot learning) ability of XNN on data from unseen dimensions. To do so, we pretrain both MPP and X-MPP on 2D PDE data and compare their finetuning performance on unseen 1D and 3D PDEs, demonstrating the benefit of multi-dimensional pretraining in XNN-based foundation models. We use three 1D PDEs and one 3D PDE: Diffusion Sorption (DS), Burgers’ equation (BG), 1D Compressible Fluid Dynamics (CFD 1D), and 3D Compressible Fluid Dynamics (CFD 3D). We evaluate using the same metric, NRMSE, and compare how finetuning performance improves as the size of the given dataset increases.
@@ 6 Conclusion
Limitations and Future Work. An axial version of the cross attention between two tensors of different dimensions is not yet fully explored. Addressing this issue as a natural extension of this work is a valuable future direction that would further widen the usability of XNNs. Additionally, improving the computational efficiency of GXNN, especially for high-dimensional data, is also a promising direction towards the practical deployment of XNNs in large-scale scientific simulations.
@@ Acknowledgement
We thank Heejun Lee for his thoughtful technical support, Jules Berman for his insightful concerns, and Kyunghyun Cho for facilitating our wonderful collaborative research. This work was partly supported by Institute of Information & communications Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT) (No.RS-2019-II190075, Artificial Intelligence Graduate School Program(KAIST); No.RS-2024-00509279, Global AI Frontier Lab; No.RS-2022-II220713, Meta-learning Applicable to Real-world Problems) and Artificial intelligence industrial convergence cluster development project funded by the Ministry of Science and ICT(MSIT, Korea) & Gwangju Metropolitan City. H. Yang was supported by the National Research Foundation of Korea(NRF) grant funded by the Korean Government(MSIT) (No. RS-2023-00279680).
@@ Appendix A Dimension-Agnostic PDE Solver
where $\operatorname{Norm}$ indicates InstanceNorm [34]. On the other hand, in the axial version, we use $\operatorname{Embedder}$ for $\psi$ in Eq.˜18 with $\max$ aggregation as follows:
$\begin{aligned} &\tilde{\boldsymbol{h}}_{d_{i}}=\max_{j\neq i,\ j\in[1,K]}\text{T}_{(K-1)(j)}\left(\operatorname{AvgPool}_{1,\ldots,K-2}\left(\operatorname{Embedder}_{K-1,K}\left(\text{T}_{(j)(K-1)(i)(K)}(\boldsymbol{x})\right)\right)\right),\\ &\boldsymbol{h}_{d_{i}}=\operatorname{Norm}(\tilde{\boldsymbol{h}}_{d_{i}}),\end{aligned}$ (25)
@@
In contrast, in the graph-based axial case, we obtain $K$ features due to the lifting layer in the patch embedding. Therefore, we also use a set-based structure but apply the same attention to each feature independently, which differs from the graph-based attention example in Eq.˜21. The self-attention used is:
@@
In the axial version, we define two $\operatorname{Deembedder}$ blocks: one for updating the node feature and the other for updating the neighborhood feature (nhbr). We aggregate them using a $\max$ operation. Then, Eq.˜22 becomes:
@@ Appendix B Proofs of Theorems
Let $\phi$ be an axis-permutation equivariant function (e.g., another SXNN or pointwise operation). The SXNN in Eq.˜9 is axis-permutation equivariant for any rank-$K$ tensor $\boldsymbol{x}\in\mathbb{R}^{d_{1}\!\times\!d_{2}\!\times\!\ldots\!\times\!d_{K}}$.
@@
We prove the axis-permutation equivariance property defined in Eq.˜6. For any axis permutation $\Pi^{\prime}\in\boldsymbol{\Pi}$, the permutation of the input in Eq.˜9 before applying $\phi$ becomes
@@
which proves that Eq.˜9 is axis-permutation equivariant for all ${\Pi^{\prime}}\in\boldsymbol{\Pi}$. ∎
@@
Under assumptions ˜1 and ˜2, the lifting layer of GXNN, Eq.˜18, is axis-permutation equivariant for any rank-$K$ (except $K=1$) tensor $\boldsymbol{x}\in\mathbb{R}^{d_{1}\!\times\!d_{2}\!\times\!\ldots\!\times\!d_{K}}$.
Proof.
The lifting layer $\operatorname{Lifting}$ described in Eq.˜18 can be equivalently written as
$\boldsymbol{h}_{d_{i}}=\phi\bigg(\text{T}_{(i)(K)}(\boldsymbol{x}),\bigoplus_{j\neq i}^{K}\text{T}_{(i)(K)}\text{T}_{(i)(K)(j)(K-1)}^{-1}\psi\left(\text{T}_{(i)(K)(j)(K-1)}(\boldsymbol{x})\right)\bigg)$ (33)
where $\text{T}_{(a)(b)(c)(d)}=\text{T}_{abcd}$ denotes the axes permutation from indices $\{1,2,3,4\}$ to $\{a,b,c,d\}$ as explained in §˜2.3. Applying an axis permutation $\Pi(\boldsymbol{x})$ to the input gives
$\phi\bigg(\text{T}_{(i)(K)}(\Pi(\boldsymbol{x})),\bigoplus_{j\neq i}^{K}\text{T}_{(i)(K)}\text{T}_{(i)(K)(j)(K-1)}^{-1}\psi\left(\text{T}_{(i)(K)(j)(K-1)}(\Pi(\boldsymbol{x}))\right)\bigg)$ (34)
Substituting $\Pi^{\prime}\text{T}_{(k)(K)}=\text{T}_{(i)(K)}\Pi$ for $i=\pi(k)$ and $\Pi^{\prime\prime}\text{T}_{(i)(K)(j)(K-1)}=\text{T}_{(k)(K)(l)(K-1)}\Pi$ for $i=\pi(k),j=\pi^{\prime}(l)$, where $\Pi^{\prime}$ permutes only axes $1$ through K1K-1 and $\Pi^{\prime\prime}$ permutes only axes $1$ through K2K-2, we get
$=\phi\bigg(\Pi^{\prime}\text{T}_{(k)(K)}(\boldsymbol{x}),\bigoplus_{\pi^{\prime}(l)\neq\pi(k)}^{K}\Pi^{\prime}\text{T}_{(k)(K)}\text{T}_{(k)(K)(l)(K-1)}^{-1}{\Pi^{\prime\prime}}^{-1}\psi\left(\Pi^{\prime\prime}\text{T}_{(k)(K)(l)(K-1)}(\boldsymbol{x})\right)\bigg).$ (35)
Let $\pi,\pi^{\prime}$ be the element-wise permutation across axes induced by $\Pi$. Since the aggregation is permutation-invariant and $\pi^{\prime}(l)\neq\pi(k)$ is equivalent to $l\neq k$ (as $\pi^{\prime}$ does not act on the $K$-th axis and $k$-th axis is transposed to $K$ by $\text{T}_{(k)(K)}$), and using ˜2 (permutation equivariance of $\psi$ under $\Pi^{\prime\prime}$), we have
$=\phi\bigg(\Pi^{\prime}\text{T}_{(k)(K)}(\boldsymbol{x}),\bigoplus_{l\neq k}^{K}\Pi^{\prime}\text{T}_{(k)(K)}\text{T}_{(k)(K)(l)(K-1)}^{-1}{\Pi^{\prime\prime}}^{-1}\Pi^{\prime\prime}\psi\left(\text{T}_{(k)(K)(l)(K-1)}(\boldsymbol{x})\right)\bigg).$ (36)
Applying ˜2 again for $\phi$ under $\Pi^{\prime}$ yields
$=\Pi^{\prime}\phi\bigg(\text{T}_{(k)(K)}(\boldsymbol{x}),\bigoplus_{l\neq k}^{K}\text{T}_{(k)(K)}\text{T}_{(k)(K)(l)(K-1)}^{-1}\psi\left(\text{T}_{(k)(K)(l)(K-1)}(\boldsymbol{x})\right)\bigg),$ (37)
which follows the definition of axes-permutation equivariance as in Eq.˜6. ∎
B.3 ˜3.3
@@
Under assumptions ˜1 and ˜3, the subsequent layers of GXNN in Eq.˜22 are axis-permutation equivariant for any rank-$K$ tensor $\boldsymbol{h}_{d_{i}}\in\mathbb{R}^{d_{1}\!\times\!d_{2}\!\times\!\ldots\!\times\!d_{K}}$.
@@
The subsequent layers $\operatorname{Subsequent}$ described in Eq.˜22 can be equivalently written as
$\boldsymbol{h}^{\prime}_{d_{i}}=\phi\left(\boldsymbol{h}_{d_{i}},\text{T}_{(i)(K)}\sum_{j=1}^{K}\text{T}_{(j)(K)}^{-1}\psi(\boldsymbol{h}_{d_{j}})\right).$ (38)
Using Eq.˜37, permutation of input leads to permutation of output of the lifting layer, which is the permutation of input of subsequent layers,
$\phi\left(\Pi^{\prime}\boldsymbol{h}_{d_{k}},\text{T}_{(\pi^{\prime}(k))(K)}\sum_{\pi^{\prime}(l)=1}^{K}\text{T}_{(\pi^{\prime}(l))(K)}^{-1}\psi(\Pi^{\prime}\boldsymbol{h}_{d_{l}})\right).$ (39)
@@
$=\Pi^{\prime}\phi\left(\boldsymbol{h}_{d_{k}},\text{T}_{(k)(K)}\sum_{l=1}^{K}\text{T}_{(l)(K)}^{-1}\psi(\boldsymbol{h}_{d_{l}})\right)=\Pi^{\prime}\boldsymbol{h}^{\prime}_{d_{k}}.$ (40)
B.4 Axes-Permutation Equivariance of Pooling Layer
Finally, the pooling layer aggregates the axis-wise features into a single feature, which remains equivariant to axis permutations $\Pi$. As described in the example in Eq.˜23, it is equivalently written as
$\boldsymbol{h}^{\prime\prime}=\sum_{j=1}^{K}\text{T}_{(j)(K)}^{-1}\boldsymbol{h}^{\prime}_{d_{j}}.$ (41)